# The Limits of Assumption-free Tests for Algorithm Performance

### v1

@@ 4.1 Background on algorithmic stability
At a high level, algorithmic stability expresses the idea that the fitted model produced by an algorithm should not change massively under a slight perturbation to the input training data. Algorithmic stability has long been a central notion in statistics and machine learning. Intuitively, it is naturally a desirable property for learning algorithms—but it also leads to many downstream implications, as well. For instance, it is known that stable algorithms tend to generalize better (Bousquet and Elisseeff,, 2002; Shalev-Shwartz et al.,, 2010). It also plays a critical role in a number of different problems, such as model selection (Meinshausen and Bühlmann,, 2010), reproducibility (Yu,, 2013), and predictive inference (Steinberger and Leeb,, 2023; Barber et al., 2021b, ). Stability is known to hold for many simple learning algorithms such as $k$-nearest neighbors (Devroye and Wagner,, 1979) and ridge regression (Bousquet and Elisseeff,, 2002). It has also been shown to hold for any algorithm whose fitted models are aggregated via bootstrapping or bagging (Breiman,, 1996; Elisseeff et al.,, 2005; Soloff et al.,, 2023).
There are also connections between stability and our ability to assess algorithmic performance; for instance, Dietterich, (1998, Section 1), in their discussion of the distinction between questions of assessing models versus assessing algorithms, pointed out that addressing questions such as EvaluateAlg must “rely on the assumption that the performance of learning algorithms changes smoothly with changes in the size of the training data”. A recent line of literature gives precise quantitative results for the problem of estimating risk under strong stability conditions: specifically, recent work by Austern and Zhou, (2020); Bayle et al., (2020) establishes asymptotic normality of the cross-validation estimator $\widehat{R}^{\textnormal{CV}}_{n}(\mathcal{A})$.
@@
Defining algorithmic stability.
@@
${\widehat{f}}_{n}=\mathcal{A}\left(\{(X_{i},Y_{i})\}_{i\in[n]};\xi\right),\quad{\widehat{f}}_{n}^{-j}=\mathcal{A}\left(\{(X_{i},Y_{i})\}_{i\in[n]\backslash\{j\}};\xi\right),$
@@ 4.2 Stability or consistency? Defining two regimes
The work of Austern and Zhou, (2020); Bayle et al., (2020), mentioned above, studied the cross-validation estimator $\widehat{R}^{\textnormal{CV}}_{n}(\mathcal{A})$ under a stability condition that is roughly equivalent to requiring
$\beta_{2}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/2}).$ (7)
@@
Each bound is a straightforward consequence of the definition of $\ell_{1}$- or $\ell_{2}$-stability. This result follows standard calculations and is similar to other results found in the stability literature, e.g., (Celisse and Guedj,, 2016, Lemma 4). (A more technical form of this result can also be established for the case of a randomized algorithm $\mathcal{A}$, but for intuition we restrict our attention to the deterministic case.)
This proposition implies that, under the strong stability assumption (7), it holds that $\mathbb{E}[(R_{P}({\widehat{f}}_{n})-R_{P,n}(\mathcal{A}))^{2}]=\mathrm{o}(1)$. Consequently, we can estimate $R_{P,n}(\mathcal{A})$ simply by estimating the risk $R_{P}({\widehat{f}}_{n})$ of a single model ${\widehat{f}}_{n}$ fitted on a training set of size $n$ (and it is relatively easy to estimate $R_{P}({\widehat{f}}_{n})$—e.g., regardless of the size of $n$, if we have $N=n+100$ then we have a holdout set of 100 data points for estimating $R_{P}({\widehat{f}}_{n})$; we do not need $N\gg n$). A similar conclusion holds if we instead assume $\beta_{1}(\mathcal{A},P,n)=\mathrm{o}(n^{-1})$.
At a high level, then, if we assume $\beta_{q}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/q})$ for either $q=1$ or $q=2$, then the distinction between questions EvaluateAlg and EvaluateModel has essentially disappeared; estimating the risk of $\mathcal{A}$ is “easy” for the trivial reason that it suffices to estimate the risk $R_{P}({\widehat{f}}_{n})$ of a single fitted model produced by $\mathcal{A}$. (Of course, we would not expect this to be the optimal strategy for estimation and inference on $R_{P,n}(\mathcal{A})$. In fact, in the setting $\beta_{2}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/2})$, the work of Austern and Zhou, (2020) establishes precise central limit theorem type result for the CV error which enables inference for $R_{P,n}(\mathcal{A})$; we do not describe the details since this is beyond the scope of the present paper.) On the other hand, in many practical settings, even relatively simple algorithms exhibit high variability and do not obey a consistency type property, i.e., $\beta_{q}(\mathcal{A},P,n)$ may be larger than $\mathcal{O}(n^{-1/q})$ or may even be $\mathcal{O}(1)$. This is well known empirically—for instance, see Soloff et al., (2023) for some examples.
We will now present our main results, showing that outside of the consistency regime there are fundamental limits on our ability to address EvaluateAlg with any black-box procedure. In particular, for $\beta_{q}(\mathcal{A},P,n)\gtrsim n^{-1/q}$, we will establish a hardness result, analogous to Theorem 2, to show that outside of the consistency regime it is impossible to perform inference on EvaluateAlg with limited data. See Figure 1 for an illustration of the phase transition between these two regimes.
@@ 4.3 Limitations under stability
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the loss $\ell$ takes values in $[0,B]$. Let ${\widehat{T}}$ be a black-box test (as in Definition 1), and assume that ${\widehat{T}}$ satisfies the stability-constrained validity condition,
$\mathbb{P}_{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)\leq\alpha\textnormal{ for all $\mathcal{A},P$ such that $\beta_{q}(\mathcal{A},P,n)\leq\gamma_{q}$ and $R_{P,n}(\mathcal{A})\geq\tau$,}$ (8)
@@
Then the power of ${\widehat{T}}$ to detect low risk is bounded as follows: for any $\mathcal{A}$ and any $P$ with $\beta_{q}(\mathcal{A},P,n)\leq\gamma_{q}$ and $R_{P,n}(\mathcal{A})<\tau$,
@@
Theorem 4 shows that, when we are outside the consistency regime (i.e., when the bound on $\ell_{q}$ stability $\gamma_{q}$ is at least as large as $\mathcal{O}(n^{-1/q})$), the problem of estimating $R_{P,n}(\mathcal{A})$ is just as hard as if we had no stability assumption at all (as in Theorem 2). Thus, there is no intermediate regime on our ability to answer the question EvaluateAlg: instead, we have a phase transition between the consistency regime, $\gamma_{q}=\mathrm{o}(n^{-1/q})$, and the hardness results of our theorems, when $\gamma_{q}\gtrsim n^{-1/q}$. As illustrated in Figure 1, we must either assume that we are in the consistency regime (in which case, it is essentially equivalent to simply answer the easier question EvaluateModel), or without this assumption, it is impossible to have high power for testing EvaluateAlg.
@@ 5.1 Problem formulation: comparing two algorithms
CompareAlg: which algorithm, $\mathcal{A}_{0}$ or $\mathcal{A}_{1}$, performs better on data drawn from $P$?
@@
CompareModel: which fitted model, ${\widehat{f}}_{0}$ or ${\widehat{f}}_{1}$, performs better on data drawn from $P$?
@@
so that $\Delta_{P}(f_{0},f_{1})$ compares the frequency of how often one function has a smaller loss than the other. From this point on, we will assume without comment that $\psi$ satisfies an antisymmetry condition,
@@
Now, given algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$ for modeling our data, we are now ready to formalize our comparison questions, CompareAlg and CompareModel.
CompareAlg: what is $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})$?  versus  CompareModel: what is $\Delta_{P}({\widehat{f}}_{0,n},{\widehat{f}}_{1,n})$?
@@ 5.2 Defining a black-box test
Consider any test ${\widehat{T}}$ that takes as input two algorithms $\mathcal{A}_{0}$, $\mathcal{A}_{1}$, a dataset $\mathcal{D}_{N}\in(\mathcal{X}\times\mathcal{Y})^{N}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D}_{N})\in\{0,1\}$. We say that ${\widehat{T}}$ is a black-box test if it can be defined as follows: for some functions $g^{(1)},g^{(2)},\dots,g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
@@
    $(\mathcal{D}^{(1)},\xi^{(1)})=g^{(1)}(\mathcal{D}_{N},\zeta^{(1)}),$
@@
    $(\mathcal{D}^{(r)},\xi^{(r)})=g^{(r)}\big{(}\mathcal{D}_{N},(\mathcal{D}^{(s)})_{1\leq s<r},({\widehat{f}}_{l}^{(s)})_{1\leq s<r;\,l=0,1},(\zeta^{(s)})_{1\leq s\leq r},(\xi^{(s)})_{1\leq s<r}\big{)},$
@@
    Repeat until some stopping criterion is reached; let ${\widehat{r}}$ denote the (data-dependent) total number of rounds.
@@
    ${\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D}_{N})=g\big{(}\mathcal{D}_{N},(\mathcal{D}^{(r)})_{1\leq r\leq{\widehat{r}}},({\widehat{f}}_{l}^{(r)})_{1\leq r\leq{\widehat{r}};\,l=0,1},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big{)}.$
@@ 5.3 Main results for algorithm comparison
In this section, we present our main result on the limit of any universally valid black-box test on the testing problem (11). Again, we would like to provide an answer to CompareAlg that is uniformly valid over all possible $\mathcal{A}_{0},\mathcal{A}_{1}$ and $P$. Specifically, fixing a error tolerance level $\alpha\in(0,1)$ and a sample size $n$, we will consider tests ${\widehat{T}}={\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D}_{N})$ that satisfy
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the comparison function $\psi$ takes values in $[-B,B]$. Let ${\widehat{T}}$ be a black-box test (as in Definition 3), and assume that ${\widehat{T}}$ satisfies the assumption-free validity condition (12).
Then the power of ${\widehat{T}}$ to compare algorithms is bounded as follows: for any $\mathcal{A}_{0},\mathcal{A}_{1}$ and any $P$ with $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})>0$,
@@
For the problem of algorithm evaluation, in the special case of a zero–one loss, we saw in Section 3.3 that the Binomial test proposed in (6) provided a concrete example of a universally valid test that achieves a power nearly as high as the upper bound given in Theorem 2. Analogously, for the problem of algorithm comparison, it is again possible to construct a similar test (in this case, for special cases such as the comparison function $\psi$ given in (10)) to show that the upper bound on power in Theorem 5 is approximately tight. Since the intuition and construction are similar to the case of algorithm comparison, we do not give the details in this work.

### v4

@@ 4.1 Background on algorithmic stability
At a high level, algorithmic stability expresses the idea that the fitted model produced by an algorithm should not change massively under a slight perturbation to the input training data. Algorithmic stability has long been a central notion in statistics and machine learning. Intuitively, it is naturally a desirable property for learning algorithms—but it also leads to many downstream implications, as well. For instance, it is known that stable algorithms tend to generalize better (Bousquet and Elisseeff,, 2002; Shalev-Shwartz et al.,, 2010). It also plays a critical role in a number of different problems, such as model selection (Meinshausen and Bühlmann,, 2010), reproducibility (Yu,, 2013), and predictive inference (Steinberger and Leeb,, 2023; Barber et al., 2021b, ). Stability is known to hold for many simple learning algorithms such as $k$-nearest neighbors (Devroye and Wagner,, 1979) and ridge regression (Bousquet and Elisseeff,, 2002). It has also been shown to hold for any algorithm whose fitted models are aggregated via bootstrapping or bagging (Breiman,, 1996; Elisseeff et al.,, 2005; Soloff et al.,, 2024).
There are also connections between stability and our ability to assess algorithmic performance; for instance, Dietterich, (1998, Section 1), in their discussion of the distinction between questions of assessing models versus assessing algorithms, pointed out that addressing questions such as EvaluateAlg must “rely on the assumption that the performance of learning algorithms changes smoothly with changes in the size of the training data”. A recent line of literature gives precise quantitative results for the problem of estimating risk under strong stability conditions: specifically, recent work by Austern and Zhou, (2020) establishes asymptotic normality of the cross-validation estimator $\widehat{R}^{\textnormal{CV}}_{n}(\mathcal{A})$. Related work by Bayle et al., (2020) also establishes asymptotic normality of $\widehat{R}^{\textnormal{CV}}_{n}(\mathcal{A})$ under weaker stability conditions.
@@
4.1.1 Defining algorithmic stability.
@@
[…]
@@ 4.2 Stability or consistency? Defining two regimes
First, we consider a regime with a strong stability assumption,
$\beta_{1}(\mathcal{A},P,n)=\mathrm{o}(n^{-1})\textnormal{ \ or \ }\beta_{2}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/2}).$ (10)
The work of Austern and Zhou, (2020), mentioned above, studies the cross-validation estimator $\widehat{R}^{\textnormal{CV}}_{n}(\mathcal{A})$ under a stability condition that is roughly equivalent to the assumption that $\beta_{2}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/2})$.  [^3] The related work of Bayle et al., (2020) places a $\mathrm{o}(n^{-1/2})$ bound on a related notion of $\ell_{2}$-stability referred to as the loss stability (also studied earlier by Kumar et al., (2013)). Loss stability is a strictly weaker stability condition, and in general is not sufficient for a consistency result (i.e., Proposition 2).
@@
Each bound is a straightforward consequence of the definition of $\ell_{1}$- or $\ell_{2}$-stability. This result follows standard calculations and is similar to other results found in the stability literature, e.g., (Celisse and Guedj,, 2016, Lemma 4). (A more technical form of this result can also be established for the case of a randomized algorithm $\mathcal{A}$, but for intuition, we restrict our attention to the deterministic case.)
Proposition 2 implies that, under the strong stability assumption (10), it holds that $\mathbb{E}[(R_{P}({\widehat{f}}_{n})-R_{P,n}(\mathcal{A}))^{2}]=\mathrm{o}(1)$. Consequently, if we assume that this condition holds, we can estimate $R_{P,n}(\mathcal{A})$ simply by estimating the risk $R_{P}({\widehat{f}}_{n})$ of a single model ${\widehat{f}}_{n}$ fitted on a training set of size $n$, which is relatively easy as long as we have a large holdout data as we have mentioned in Section 3.1.
At a high level, then, if we assume $\beta_{q}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/q})$ for either $q=1$ or $q=2$, then the distinction between questions EvaluateAlg and EvaluateModel has essentially disappeared; estimating the risk of $\mathcal{A}$ is “easy” for the trivial reason that it suffices to estimate the risk $R_{P}({\widehat{f}}_{n})$ of a single fitted model produced by $\mathcal{A}$. (Of course, we would not expect this to be the optimal strategy for estimation and inference on $R_{P,n}(\mathcal{A})$. In fact, in the setting $\beta_{2}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/2})$ along with some additional assumptions, the work of Austern and Zhou, (2020) establishes precise central limit theorem type result for the CV error which enables inference for the algorithm risk; we do not describe the details since this is beyond the scope of the present paper.) A few common algorithms which satisfy this level of stability are $\ell_{q}$ regularized algorithms ($q>1$) or soft-margin SVM (Bousquet and Elisseeff,, 2002; Elisseeff et al.,, 2005; Wibisono et al.,, 2009). On the other hand, in many practical settings, even relatively simple algorithms exhibit high variability and do not obey a consistency type property, i.e., $\beta_{q}(\mathcal{A},P,n)$ may be larger than $\mathcal{O}(n^{-1/q})$ or may even be $\mathcal{O}(1)$. This happens for instance in linear regression when $d\approx n$ or in some sparse algorithms (Xu et al.,, 2011).
We will now present our main results, showing that outside of the consistency regime, there are fundamental limits on our ability to address EvaluateAlg with any black-box procedure. In particular, for $\beta_{q}(\mathcal{A},P,n)\gtrsim n^{-1/q}$, we will establish a hardness result, analogous to Theorem 1, to show that outside of the consistency regime it is impossible to perform inference on EvaluateAlg with limited data. See Figure 1 for an illustration of the phase transition between these two regimes.
@@ 4.3 Limitations under stability
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the loss $\ell$ takes values in $[0,B]$. Let $^{\widehat{T}}$ be a black-box test (as in Definition 1), and assume that $^{\widehat{T}}$ satisfies the stability-constrained validity condition,
[…]
@@
Then the power of $^{\widehat{T}}$ to detect low risk is bounded as follows: for any $\mathcal{A}$ and any $P$ with $\beta_{q}(\mathcal{A},P,n)\leq\gamma_{q}$ and $R_{P,n}(\mathcal{A})<\tau$,
@@
Theorem 4 shows that, when we are outside the consistency regime (i.e., when the bound on $\ell_{q}$ stability $\gamma_{q}$ is at least as large as $\mathcal{O}(n^{-1/q})$), the problem of estimating $R_{P,n}(\mathcal{A})$ is just as hard as if we had no stability assumption at all (as in Theorem 1). Thus, there is no intermediate regime on our ability to answer the question EvaluateAlg: instead, we have a phase transition between the consistency regime, $\gamma_{q}=\mathrm{o}(n^{-1/q})$, and the hardness results of our theorems, when $\gamma_{q}\gtrsim n^{-1/q}$. As illustrated in Figure 1, we must either assume that we are in the consistency regime (in which case, it is essentially equivalent to simply answer the easier question EvaluateModel), or without this assumption, it is impossible to have high power for testing EvaluateAlg. Theorem 4 also reveals that the the stability assumption $\beta_{2}(\mathcal{A},P,n)=\mathrm{o}(n^{-1/2})$ assumed in Austern and Zhou, (2020) for inference of $R_{P,n}(\mathcal{A})$ via CV error is necessary in general.
Remark 4.
As we have mentioned before, our results are built on the black-box testing framework established in Kim and Barber, (2023). The focus of their work (i.e., establishing the hardness of testing algorithmic stability) is different than the aim of inference on algorithmic risk considered here; however, some of the proof techniques are related, particularly in the constructions of counterexamples. Here we briefly comment on some of the ways that this work extends the technical tools developed by Kim and Barber, (2023). First, our work allows for more general X and Y settings (see some discussions in Remark 3 and Appendix B.1) and and it paves the way for considering the limits of tests with computational budget constraints in settings when $|\mathcal{X}|$ and $|\mathcal{Y}|$ are finite. Partial progress has been made in Luo and Barber, (2024). Second, compared to the existing framework, we also considered relaxing the assumption-free validity in studying EvaluateAlg by exploring how an algorithm stability assumption helps in EvaluateAlg. The proof of the negative result in Theorem 4 also goes beyond the original framework in Kim and Barber, (2023).
@@ 5.1 Problem formulation: comparing two algorithms
[…]
@@
[…]
@@
so that $\Delta_{P}(f_{0},f_{1})$ compares the frequency of how often one function has a smaller loss than the other. From this point on, we will assume implicitly that $\psi$ satisfies an antisymmetry condition,
@@
Now, given algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$ for modeling our data, we are ready to formalize our comparison questions, CompareAlg and CompareModel.
@@
The question CompareModel compares the fitted models ${\widehat{f}}_{0,n},{\widehat{f}}_{1,n}$ obtained by training these algorithms on a particular data set of size $n$, given by ${\widehat{f}}_{l,n}=\mathcal{A}_{l}\big(\{(X_{i},Y_{i})\}_{i\in[n]};\xi\big)$ for each l=0,1l=0,1.444In this section we take the convention that the two algorithms, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$, are run with the same random seed $\xi$. The same types of results will hold regardless of this choice, with appropriate modifications to the definitions.  In contrast, the question CompareAlg takes an expected value over the training set,
@@ 5.2 Defining a black-box test
Let $\mathfrak{D}=\cup_{m\geq 0}(\mathcal{X}\times\mathcal{Y})^{m}$ denotes the space of data sets of any size. Consider any test $^{\widehat{T}}$ that takes as input two algorithms $\mathcal{A}_{0}$, $\mathcal{A}_{1}$, a dataset $\mathcal{D}\in\mathfrak{D}$, and returns a (possibly randomized) output ${\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D})\in\{0,1\}$. We say that $^{\widehat{T}}$ is a black-box test if it can be defined as follows: for some functions $g^{(1)},g^{(2)},\dots,g$, and for random seeds, $\zeta^{(1)},\zeta^{(2)},\dots,\zeta\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\textnormal{Unif}[0,1]$,
@@
    $(\mathcal{D}^{(1)},\xi^{(1)})=g^{(1)}(\mathcal{D},\zeta^{(1)}),$
@@
    $(\mathcal{D}^{(r)},\xi^{(r)})=g^{(r)}\big(\mathcal{D},(\mathcal{D}^{(s)})_{1\leq s<r},({\widehat{f}}_{l}^{(s)})_{1\leq s<r;\,l=0,1},(\zeta^{(s)})_{1\leq s\leq r},(\xi^{(s)})_{1\leq s<r}\big),$
@@
    Repeat until some stopping criterion is reached; let $^{\widehat{r}}$ denote the (data-dependent) total number of rounds.
@@
    ${\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D})=g\big(\mathcal{D},(\mathcal{D}^{(r)})_{1\leq r\leq{\widehat{r}}},({\widehat{f}}_{l}^{(r)})_{1\leq r\leq{\widehat{r}};\,l=0,1},(\zeta^{(r)})_{1\leq r\leq{\widehat{r}}},(\xi^{(r)})_{1\leq r\leq{\widehat{r}}},\zeta\big).$
@@ 5.3 Main results for algorithm comparison
In this section, we present our main result on the limit of any universally valid black-box test on the testing problem (14). Again, we would like to provide an answer to CompareAlg that is uniformly valid over all possible $\mathcal{A}_{0},\mathcal{A}_{1}$ and $P$. Specifically, fixing a error tolerance level $\alpha\in(0,1)$ and an evaluation sample size $n$, an available dataset $\mathcal{D}_{N}=\{(X_{i},Y_{i})\}_{i\in[N]}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$ of size $N$, we will consider tests ${\widehat{T}}={\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D}_{N})$ that satisfy
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the comparison function $\psi$ takes values in [B,B][-B,B]. Let $^{\widehat{T}}$ be a black-box test (as in Definition 3), and assume that $^{\widehat{T}}$ satisfies the assumption-free validity condition (15).
Then the power of $^{\widehat{T}}$ to compare algorithms is bounded as follows: for any $\mathcal{A}_{0},\mathcal{A}_{1}$ and any $P$ with $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})>0$,
@@
For the problem of algorithm evaluation, in the special case of a zero-one loss, we saw in Section 3.2 that the Binomial test proposed in (9) provided a concrete example of a universally valid test that achieves a power nearly as high as the upper bound given in Theorem 1. Analogously, for the problem of algorithm comparison, it is again possible to construct a similar test (in this case, for special cases such as the comparison function $\psi$ given in (13)) to show that the upper bound on power in Theorem 5 is approximately tight. Since the intuition and construction are similar to the case of algorithm comparison, we do not give the details in this work.